OpenStudy (anonymous):

Which function has the greatest rate of change? A. 6x – 7y = 21 B. y = 2x – 2 C. x y 2 7 3 13 4 19 D. A linear function that goes through point 1, 2 and point 0, negative 1

OpenStudy (amistre64):

since they are all linear, define the slopes

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

i mean, for each one you will want to define the slope in order to compare them. the slope tells us about the rate of change

OpenStudy (anonymous):

so I have to find what the slope is for each one

OpenStudy (amistre64):

correct.

OpenStudy (amistre64):

what do you know about finding slopes?

OpenStudy (anonymous):

ok I have to refresh quick

OpenStudy (anonymous):

y2-y1 ------ x2-x1

OpenStudy (amistre64):

slope formula is good when we are given point to play with, yes

OpenStudy (amistre64):

so for the last 2, we could use the slope formula. and for the first 2, we can use the form of the equation: y = mx + b, where m is the value of the slope

OpenStudy (amistre64):

A. 6x – 7y = 21 solving for y we get: y = 6/7 x - 3 ^^^ B. y = 2x – 2 ^ C. x y 2 7 3 13 4 19 13-7 ---- 3-2 D. A linear function that goes through point 1, 2 and point 0, negative 1 2--1 ----- 1-0

OpenStudy (amistre64):

refresh the page to get rid of any odd diamond marks

OpenStudy (anonymous):

its good now

OpenStudy (anonymous):

C 6/1 D 1/1 which would be 1

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

C = 6 and D = 3, recall that subtracting a negative is adding.

OpenStudy (amistre64):

so, 6/7 is less than 2, so A is out 2 is less than 6 so B is out 6 is bigger than 3, so D is out :)

OpenStudy (anonymous):

so C

OpenStudy (amistre64):

C is my pick. 6 is the biggest rate of change

OpenStudy (anonymous):

thank you